OpenStudy (anonymous):

find -f(x) when f(x)=-4+1

OpenStudy (mathmate):

f(x)=-4+1=-3 so -f(x)=-(-3)=3

OpenStudy (anonymous):

You just take the opposite of everything

OpenStudy (mathmate):

unless f(x)=-4x+1, then -f(x)=-(-4x+1)=4x-1

OpenStudy (anonymous):

thank u..

OpenStudy (mathmate):

yw

OpenStudy (uri):

@cralls18 Welcome to open study xD

OpenStudy (anonymous):

thanks uri

OpenStudy (uri):

yw :D